OpenStudy (anonymous):

rewrite log5-log200 with a single log then solve

OpenStudy (anonymous):

log(5/200) = log(1/40) = -1.60205999

OpenStudy (anonymous):

thanks!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

the difference of like logs is equal to a log of their division (order sensitive!)

OpenStudy (anonymous):

ok thanks dude

OpenStudy (anonymous):

loga-logb = log(a/b)

OpenStudy (anonymous):

loga+logb = log(ab)

OpenStudy (anonymous):

SAME base, order sensitive in division

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

how about......... 5logx-4log(x-3)

OpenStudy (anonymous):

you better give me a medal :P

OpenStudy (anonymous):

how do i give it to you?

OpenStudy (anonymous):

ill give as many as i can though,..lol

OpenStudy (anonymous):

= logx^5-log(x-3)^4 = log((x^5)/((x-3)^4)

OpenStudy (anonymous):

how can i give you more medals? =P

OpenStudy (anonymous):

coefficients are the same as exponents because logs are power functions, so that log(base a)b=c is equivalent to a^c=b log means what i raise the base to to get the number "what do i raise a to to get b" = c

OpenStudy (anonymous):

new medal = new thread

OpenStudy (anonymous):

so 1 medal per thing you type?

OpenStudy (anonymous):

factor 9x^4 - y^6

OpenStudy (anonymous):

ask new questions in new threads :D

OpenStudy (anonymous):

ok answer them for me???

OpenStudy (anonymous):

how do you ask a new question?